# VLM2Vec: Training Vision-Language Models for Massive Multimodal Embedding Tasks

### v2

@@ 4.4.4 Impact of Instructions
Previous studies have shown the influence of instructions on addressing various tasks. Vlm2Vec, which leverages a VLM as its backbone and is trained on large-scale datasets with instructions, is expected to better generalize across tasks and improve performance in multimodal embedding tasks. In this section, we evaluate the performance of Vlm2Vec with and without task-specific instructions to quantify the impact of incorporating instructions into the embedding process. As shown in Table 4, excluding instructions leads to an average performance drop of around 30%, highlighting the importance of instruction-guided embeddings.
Table 4: Comparison of Vlm2Vec with and without task-specific instructions. The percentage decrease highlights the significant performance drop when instructions are not incorporated.
@@ 5.1 Text Embedding
Text embeddings have demonstrated significant potential in powering downstream applications such as information retrieval (Karpukhin et al., 2020; Xiong et al., 2020), text similarity (Gao et al., 2021b), prompt retrieval for in-context learning (Hongjin et al., 2022), and classification (Logeswaran & Lee, 2018; Reimers & Gurevych, 2019). Early work focused on creating effective embeddings for specific tasks. With the rise of pretrained language models, efforts have shifted toward developing universal embedding models capable of handling a wide range of embedding tasks. Studies such as GTR (Ni et al., 2022) and E5 (Wang et al., 2022a) leveraged large amounts of noisy paired data to pretrain and fine-tune dense retrievers. More recent works like TART (Asai et al., 2022) and InstructOR (Su et al., 2023) introduced natural language prompts to guide embedding models in producing task-relevant embeddings. Building on this, models like E5Mistral(Wang et al., 2024), SFR-Embedding(Meng et al., 2024), RepLLaMA(Ma et al., 2024b), GTE-Qwen2(Li et al., 2023b), and NV-Embed (Lee et al., 2024) have utilized pretrained large language models (LLMs) as their backbone, fine-tuning them with multi-task data and instructions. These models have delivered significant improvements over earlier approaches that did not use LLMs for initialization or instruction tuning. However, these advances come at a cost: both the model size and the dimensionality of output vectors have increased substantially, leading to higher costs.
@@ 5.2 Multimodal Embeddings
Most research on universal multimodal embeddings involves fine-tuning models like CLIP or BLIP, typically using simple fusion mechanisms to combine visual and language information. For instance, UniIR (Wei et al., 2023) creates multimodal embeddings by simply adding text and visual features, while MagicLens (Zhang et al., 2024) employs shallow self-attention layers to integrate these features more effectively. The study most similar to ours is E5-V (Jiang et al., 2024), which converts a multimodal generative model into an embedding model. We compare our approach with E5-V and find that our models achieve significant improvements in performance.
@@ 5.3 Embedding Benchmarks
Significant efforts have been made to develop benchmarks for evaluating retrieval systems. For text retrieval models, MS MARCO (Nguyen et al., 2016) and Natural Questions (Kwiatkowski et al., 2019b) are two of the most widely used benchmarks in general domains. To broaden the evaluation across more diverse domains, BEIR (Thakur et al., ) was introduced, incorporating 18 datasets from various fields. Building on this, MTEB (Muennighoff et al., 2023) further expands BEIR’s scope by adding more tasks, such as classification, clustering, and semantic textual similarity (STS), to assess the generalization capabilities of embedding models.
For multimodal retrieval, several benchmarks have been introduced to evaluate model performance across different modalities. MBEIR (Wei et al., 2023) includes 8 tasks and 16 datasets, designed to test models’ ability to retrieve information based on various forms of queries (text, image, or a combination) and instructions that span multiple modalities.
@@ 6 Conclusion
Appendix A Appendix
@@ A.1.4 Visual Grounding
Table 5: The detailed results of the baselines and our Vlm2Vec on MMEB, which includes 20 in-distribution datasets and 16 out-of-distribution datasets. The out-of-distribution datasets are highlighted with a yellow background in the table. We only include the best version of Vlm2Vec in the table, which was trained using LoRA.
@@
ImageNet-1K 55.8 63.5 45.4 10.3 48.0 9.6 53.7 65.6
N24News 34.7 38.6 13.9 36.0 33.7 23.4 33.9 79.5
HatefulMemes 51.1 51.7 47.2 49.6 49.0 49.7 51.0 67.1
VOC2007 50.7 52.4 64.3 52.1 51.6 49.9 62.7 88.6
SUN397 43.4 68.8 39.6 34.5 57.0 33.1 61.7 72.7
Place365 28.5 37.8 20.0 21.5 31.5 8.6 38.0 42.6
ImageNet-A 25.5 14.2 42.6 3.2 8.0 2.0 12.9 19.3
ImageNet-R 75.6 83.0 75.0 39.7 70.9 30.8 61.6 70.2
ObjectNet 43.4 51.4 40.3 20.6 31.6 7.5 37.1 29.5
Country-211 19.2 16.8 14.2 2.5 6.2 3.1 8.8 13.0
All Classification 42.8 47.8 40.3 27.0 38.8 21.8 42.1 54.8
@@
OK-VQA 7.5 11.5 2.4 8.7 12.7 8.9 24.5 63.2
A-OKVQA 3.8 3.3 1.5 3.2 2.9 5.9 10.6 50.2
DocVQA 4.0 5.3 4.2 2.6 3.0 1.7 5.6 78.4
InfographicsVQA 4.6 4.6 2.7 2.0 5.9 2.3 5.0 40.8
ChartQA 1.4 1.5 3.0 0.5 0.9 2.4 1.8 59.0
Visual7W 4.0 2.6 1.2 1.3 2.5 5.8 12.3 47.7
ScienceQA 9.4 10.2 7.9 6.8 5.2 3.6 11.6 43.4
VizWiz 8.2 6.6 2.3 4.0 1.7 2.6 19.2 39.2
GQA 41.3 52.5 57.5 9.7 43.5 7.8 49.3 60.7
TextVQA 7.0 10.9 1.0 3.3 4.6 8.2 10.3 66.1
All VQA 9.1 10.9 8.4 4.2 8.3 4.9 15.0 54.9
@@
VisDial 30.7 25.4 21.5 18.0 24.8 9.2 37.6 73.3
CIRR 12.6 15.4 15.1 9.8 39.1 6.1 53.2 47.8
VisualNews_t2i 78.9 74.0 51.0 48.1 50.7 13.5 63.6 67.2
VisualNews_i2t 79.6 78.0 52.4 13.5 21.1 8.1 68.8 70.7
MSCOCO_t2i 59.5 63.6 58.3 53.7 54.1 20.7 72.0 70.6
MSCOCO_i2t 57.7 62.1 55.0 20.3 40.0 14.0 74.1 66.5
NIGHTS 60.4 66.1 62.9 56.5 58.1 4.2 69.7 66.1
WebQA 67.5 62.1 58.1 55.4 43.0 17.7 86.3 88.1
FashionIQ 11.4 13.8 20.1 9.3 11.2 2.8 39.3 12.9
Wiki-SS-NQ 55.0 44.6 55.1 28.7 18.7 8.6 11.3 56.6
OVEN 41.1 45.0 56.0 39.5 1.6 5.9 66.6 47.3
EDIS 81.0 77.5 23.6 54.4 62.6 26.8 78.2 79.9
All Retrieval 53.0 52.3 31.6 33.9 35.4 11.5 60.1 62.3
@@
MSCOCO 33.8 34.5 46.4 28.9 22.1 10.8 46.6 67.3
RefCOCO 56.9 54.2 70.8 47.4 22.8 11.9 67.8 84.7
RefCOCO-matching 61.3 68.3 50.8 59.5 35.6 38.9 62.9 79.2
Visual7W-pointing 55.1 56.3 70.1 52.0 23.4 14.3 71.3 86.8
All Visual Grounding 51.8 53.3 59.5 47.0 26.0 19.0 62.2 79.5

### v3

@@ 4.3.4 Impact of Instructions
Previous studies have shown the influence of instructions on addressing various tasks. Vlm2Vec, which leverages a VLM as its backbone and is trained on large-scale datasets with instructions, is expected to better generalize across tasks and improve performance in multimodal embedding tasks. In this section, we evaluate the performance of both CLIP and Vlm2Vec with and without task-specific instructions to quantify the impact of incorporating instructions into the embedding process. As shown in Table 4, incorporating instructions reduces the CLIP model’s performance by 29.4%, while our Vlm2Vec achieves a 49.4% improvement. This highlights how a VLM backbone enhances the embedding model’s instruction-following capability and emphasizes the advantages of instruction-guided embeddings.
Table 4: Comparison of CLIP and our Vlm2Vec with and without task-specific instructions. Incorporating instructions could decrease CLIP’s performance by 29.4%, whereas our Vlm2Vec achieves a 49.4% improvement. Vlm2Vec utilizes Phi-3.5-V as its backbone.
@@
CILP
w/o instruction 42.8 9.1 53.0 51.8 37.1 38.7 37.8
w/ instruction 17.4 8.0 41.3 52.9 23.8 30.3 26.7
$\Delta$59.3% -12.1% -22.1% 2.1% -35.8% -21.7% -29.4%
Ours (Vlm2Vec)
w/o instruction 36.7 33.5 31.1 44.3 37.3 31.6 34.8
@@ 5.1 Text Embedding
Text embeddings have demonstrated significant potential in powering downstream applications such as information retrieval (Karpukhin et al., 2020; Xiong et al., 2020), text similarity (Gao et al., 2021b), prompt retrieval for in-context learning (Hongjin et al., 2022), and classification (Logeswaran & Lee, 2018; Reimers & Gurevych, 2019). Early work focused on creating effective embeddings for specific tasks. With the rise of pretrained language models, efforts have shifted toward developing universal embedding models capable of handling a wide range of embedding tasks. Studies such as GTR (Ni et al., 2022) and E5 (Wang et al., 2022a) leveraged large amounts of noisy paired data to pretrain and fine-tune dense retrievers. More recent works like TART (Asai et al., 2022) and InstructOR (Su et al., 2023) introduced natural language prompts to guide embedding models in producing task-relevant embeddings. Building on this, models like E5Mistral(Wang et al., 2024), SFR-Embedding(Meng et al., 2024), RepLLaMA(Ma et al., 2024b), GTE-Qwen2(Li et al., 2023b), and NV-Embed (Lee et al., 2024) have utilized pretrained large language models (LLMs) as their backbone, fine-tuning them with multi-task data and instructions. These models have delivered significant improvements over earlier approaches that did not use LLMs for initialization or instruction tuning.
@@ 5.2 Multimodal Embeddings
Most research on universal multimodal embeddings involves fine-tuning models like CLIP or BLIP, typically using simple fusion mechanisms to combine visual and language information. For instance, UniIR (Wei et al., 2023) creates multimodal embeddings by simply adding text and visual features, while MagicLens (Zhang et al., 2024) employs shallow self-attention layers to integrate these features more effectively. The study most similar to ours is E5-V (Jiang et al., 2024), a contemporary work that fine-tunes a vision-language model using only text training data.
@@ 5.3 Embedding Benchmarks
Significant efforts have been made to develop benchmarks for evaluating retrieval systems. For text retrieval models, MS MARCO (Nguyen et al., 2016) and Natural Questions (Kwiatkowski et al., 2019b) are two of the most widely used benchmarks in general domains. To broaden the evaluation across more diverse domains, BEIR (Thakur et al., ) was introduced, incorporating 18 datasets from various fields. Building on this, MTEB (Muennighoff et al., 2023) further expands BEIR’s scope by adding more tasks, such as classification, clustering, and semantic textual similarity (STS).
For multimodal retrieval, several benchmarks have been introduced to evaluate model performance across different modalities. MBEIR (Wei et al., 2023) includes 8 tasks and 16 datasets, designed to test models’ ability to retrieve information based on various forms of queries and instructions.
@@ 6 Conclusion
Appendix A Details of MMEB
In this section, we provide additional details about our proposed benchmark, MMEB (Massive Multimodal Embedding Benchmark). Section A.1 outlines the specifics of the 36 datasets used in the MMEB benchmark. Section A.2 explains the process for determining the number of candidates in MMEB.
@@ A.1.4 Visual Grounding
A.2 Selection of Number of Candidates
A large number of candidates can make the benchmark more challenging and realistic. However, we also considered the computational cost when designing the benchmark. Choosing an excessively large number of candidates could result in very high inference costs, which may hinder rapid model iteration. As shown in Table 5, we compare the performance of Vlm2Vec with different numbers of candidates in the MMEB benchmark. The results show that if the number of candidates is too small, the benchmark becomes saturated quickly. To balance evaluation cost with benchmark difficulty, we selected 1,000 as the optimal number of candidates.
Table 5: We compare the performance of Vlm2Vec using different numbers of candidates in MMEB. To balance evaluation cost with benchmark difficulty, we selected 1,000 as the optimal number of candidates.
#Candidates Meta-Task Average Score Average Score
Classification VQA Retrieval Grounding IND OOD Overall
# of datasets $\rightarrow$ 10 10 12 4 20 16 36
100 54.8 81.8 86.1 89.6 85.2 65.9 76.6
500 54.8 65.9 72.6 82.8 74.6 57.3 66.9
1000 54.8 54.9 62.3 79.5 66.5 52.0 60.1
2000 54.8 50.1 56.7 71.0 62.2 48.0 55.9
5000 54.8 41.3 46.5 65.3 54.5 43.2 49.5
Table 6: The detailed results of the baselines and our Vlm2Vec on MMEB, which includes 20 in-distribution datasets and 16 out-of-distribution datasets. The out-of-distribution datasets are highlighted with a yellow background in the table. We include only the best version of Vlm2Vec in the table, which uses LLaVA-1.6 as backbone.
@@ A.2 Selection of Number of Candidates
ImageNet-1K 55.8 63.5 45.4 10.3 48.0 9.6 58.3 74.5
N24News 34.7 38.6 13.9 36.0 33.7 23.4 42.5 80.3
HatefulMemes 51.1 51.7 47.2 49.6 49.0 49.7 56.4 67.9
VOC2007 50.7 52.4 64.3 52.1 51.6 49.9 66.2 91.5
SUN397 43.4 68.8 39.6 34.5 57.0 33.1 63.2 75.8
Place365 28.5 37.8 20.0 21.5 31.5 8.6 36.5 44.0
ImageNet-A 25.5 14.2 42.6 3.2 8.0 2.0 9.8 43.6
ImageNet-R 75.6 83.0 75.0 39.7 70.9 30.8 66.2 79.8
ObjectNet 43.4 51.4 40.3 20.6 31.6 7.5 32.2 39.6
Country-211 19.2 16.8 14.2 2.5 6.2 3.1 11.3 14.7
All Classification 42.8 47.8 40.3 27.0 38.8 21.8 44.3 61.2
@@
OK-VQA 7.5 11.5 2.4 8.7 12.7 8.9 25.4 69.0
A-OKVQA 3.8 3.3 1.5 3.2 2.9 5.9 8.8 54.4
DocVQA 4.0 5.3 4.2 2.6 3.0 1.7 6.2 52.0
InfographicsVQA 4.6 4.6 2.7 2.0 5.9 2.3 4.6 30.7
ChartQA 1.4 1.5 3.0 0.5 0.9 2.4 1.6 34.8
Visual7W 4.0 2.6 1.2 1.3 2.5 5.8 14.5 49.8
ScienceQA 9.4 10.2 7.9 6.8 5.2 3.6 12.8 42.1
VizWiz 8.2 6.6 2.3 4.0 1.7 2.6 24.3 43.0
GQA 41.3 52.5 57.5 9.7 43.5 7.8 48.8 61.2
TextVQA 7.0 10.9 1.0 3.3 4.6 8.2 15.1 62.0
All VQA 9.1 10.9 8.4 4.2 8.3 4.9 16.2 49.9
@@
VisDial 30.7 25.4 21.5 18.0 24.8 9.2 42.2 80.9
CIRR 12.6 15.4 15.1 9.8 39.1 6.1 51.3 49.9
VisualNews_t2i 78.9 74.0 51.0 48.1 50.7 13.5 74.3 75.4
VisualNews_i2t 79.6 78.0 52.4 13.5 21.1 8.1 76.8 80.0
MSCOCO_t2i 59.5 63.6 58.3 53.7 54.1 20.7 68.5 75.7
MSCOCO_i2t 57.7 62.1 55.0 20.3 40.0 14.0 72.1 73.1
NIGHTS 60.4 66.1 62.9 56.5 58.1 4.2 66.2 65.5
WebQA 67.5 62.1 58.1 55.4 43.0 17.7 89.6 87.6
FashionIQ 11.4 13.8 20.1 9.3 11.2 2.8 40.2 16.2
Wiki-SS-NQ 55.0 44.6 55.1 28.7 18.7 8.6 12.2 60.2
OVEN 41.1 45.0 56.0 39.5 1.6 5.9 69.4 56.5
EDIS 81.0 77.5 23.6 54.4 62.6 26.8 79.2 87.8
All Retrieval 53.0 52.3 31.6 33.9 35.4 11.5 61.8 67.4
@@
MSCOCO 33.8 34.5 46.4 28.9 22.1 10.8 46.6 80.6
RefCOCO 56.9 54.2 70.8 47.4 22.8 11.9 67.8 88.7
RefCOCO-matching 61.3 68.3 50.8 59.5 35.6 38.9 62.9 84.0
Visual7W-pointing 55.1 56.3 70.1 52.0 23.4 14.3 71.3 90.9
All Visual Grounding 51.8 53.3 59.5 47.0 26.0 19.0 65.3 86.1
@@
Table 11: Zero-shot text-image retrieval performance on Flickr30K. As a general multimodal representation model, Vlm2Vec can still achieve competitive T2I (Text-to-Image) and I2T (Image-to-Text) scores when compared to existing CLIP-like models. The baseline numbers are sourced from Sun et al. (2023) and  Zhang et al. (2024). We use the best version of Vlm2Vec here, which is built upon the LLaVA-1.6 backbone.
Model image retrieval text retrieval
R@1 R@5 R@10 R@1 R@5 R@10
OpenAI CLIP-B/16 62.1 85.6 91.8 81.9 96.2 98.8
Open CLIP-B/16 69.8 90.4 94.6 86.3 97.9 99.4
EVA-02-CLIP-B/16 71.2 91.0 94.7 85.7 96.7 98.9
OpenAI CLIP-L/14 65.2 87.3 92.0 85.2 97.3 99.0
Open CLIP-L/14 75.0 92.5 95.6 88.7 98.4 99.2
EVA-02-CLIP-L/14 77.3 93.6 96.8 89.7 98.6 99.2
MagicLens-B 76.2 93.7 96.5 87.9 97.7 99.5
MagicLens-L 79.7 95.0 97.4 89.6 98.7 99.4
Vlm2Vec 80.3 95.0 97.4 94.6 99.5 99.8